### v1

@@ Abstract
Binary classification tasks with imbalanced classes pose significant challenges in machine learning. Traditional classifiers often struggle to accurately capture the characteristics of the minority class, resulting in biased models with subpar predictive performance. In this paper, we introduce a novel approach to tackle this issue by leveraging Universum points to support the minority class within quadratic twin support vector machine models. Unlike traditional classifiers, our models utilize quadratic surfaces instead of hyperplanes for binary classification, providing greater flexibility in modeling complex decision boundaries. By incorporating Universum points, our approach enhances classification accuracy and generalization performance on imbalanced datasets. We generated four artificial datasets to demonstrate the flexibility of the proposed methods. Additionally, we validated the effectiveness of our approach through empirical evaluations on benchmark datasets, showing superior performance compared to conventional classifiers and existing methods for imbalanced classification.
Keywords:
Binary classification, Quadratic support vector machine, Twin support vector machine, Class imbalance, Universum data.
@@ 1 Introduction
However, many real-world applications pose challenges as they often involve data that is not linearly separable. Addressing this requires robust techniques capable of handling nonlinear situations effectively. One well-established approach is to leverage the concept of kernel methods, which relies on the existence of a nonlinear mapping. This mapping transforms the original data into a higher-dimensional, potentially infinite-dimensional, feature space where the data becomes linearly separable. Notably, the appeal of kernel methods lies in their ability to bypass the need for explicitly knowing this mapping. Instead, they focus on selecting appropriate kernel functions with advantageous properties. However, the challenge lies in determining the suitable kernel function, and fine-tuning the associated hyperparameters can be computationally demanding. As a result, there is a growing interest in practical methods that seek nonlinear classifiers directly in the original feature space.
Kernel-free models offer a direct approach for handling nonlinearly separable datasets without the need to map them to a higher-dimensional feature space. The quadratic surface support vector machine (QSSVM) dagher2008quadratic utilizes quadratic surfaces for class separation. Bai et al.Bai2015 introduced a kernel-free least square QSSVM tailored for disease diagnosis. A kernel-free support vector machine, utilizing a double-well potential, proposed in gao2021kernel , aims to achieve a specialized fourth-order polynomial separating surface. However, when pursuing nonlinear classifiers, the complexity of the optimization program significantly increases due to the involvement of numerous decision variables, leading to computational challenges and potential overfitting. To mitigate this, incorporating a surrogate that promotes sparsity among decision variables proves advantageous moosaei2023sparse ; mousavi2019quadratic ; mousavi2020survey . To achieve less expensive quadratic programs, gao2019quadratic discusses a least squares twin QSSVM, leveraging two quadratic surfaces for data representation.
@@
Our methodology draws inspiration from the approach introduced in richhariya2020reduced , which incorporates Universum points to bolster the minority class in linearly separable datasets. Building upon this foundation, we extend its application to the framework of quadratic twin support vector machines (QTSVM). Leveraging the inherent flexibility of QTSVM, which employs quadratic surfaces instead of hyperplanes for binary classification, we aim to enhance the efficacy of Universum points in addressing class imbalance. This innovative integration enables us to capture intricate class relationships and adapt to complex data distributions more effectively. In addition to the hinge loss, we derive the least square formulation, which adopts the $\ell_{2}$ norm to penalize slack variables and replaces inequality and equality constraints. To validate the effectiveness of our proposed methods, we conduct experiments on various datasets, including artificial and public benchmark datasets. These experiments demonstrate the efficacy and efficiency of our proposed models compared to other well-known SVM models.
@@
[…]
@@ 2 Problem Statement and Related Works
2.1 Support Vector Machine
Suppose that we are faced with a binary classification problem and the related data set is $T=\left\{(x_{1},y_{1}),\dots,(x_{m},y_{m})\right\}\in(\mathbb{R}^{n}\times\{\pm 1\})^{m},$ where the $x_{i}$s are $n$-dimensional samples and the $y_{i}$s are their corresponding labels. For such a two-class problem when the classes are (almost) linearly separable, the SVM algorithm finds two parallel hyperplanes with the maximal minimum margin. The training samples from the two classes are divided by the middle hyperplane $f(x)=w^{T}x+b=0$ if and only if $y_{i}(w^{T}x_{i}+b)\geq 1,\ i=1,\ldots,m.$ The hyperplane for the almost separable case can be obtained by solving the following optimization problem:
$\mathop{\min}\limits_{w,b,\xi}\,$ $\frac{1}{2}{{\left\|w\right\|}^{2}}+\mu\sum_{i=1}^{m}\xi_{i}$
s.t. $y_{i}(x^{T}_{i}w+b)\geq 1-\xi_{i},\quad i=1,\dots,m,$
$w\in\mathbb{R}^{n},\,b\in\mathbb{R},\ \xi\in\mathbb{R}^{m}_{+},$
where $\mu>0$ is a penalty parameter to control the trade-off and $\xi_{i}$’s are slack variables to accommodate noisy data points and outliers. The above soft margin SVM exhibits several desirable properties, including robustness to noise and outliers, as it minimizes the effect of individual misclassified points while maximizing the margin. However, addressing the dual of (2.1) is computationally advantageous due to its more structured nature. The Lagrangian dual problem is as follows:
$\mathop{\min_{\alpha}}$ $\frac{1}{2}\sum\limits_{i=1}^{m}\sum\limits_{j=1}^{m}\alpha_{i}\alpha_{j}y_{i}y_{j}x_{i}^{T}x_{j}-\sum\limits_{i=1}^{m}\alpha_{i}$
s.t. $\sum\limits_{i=1}^{m}y_{i}\alpha_{i}=0,$
$0\leq\alpha_{i}\leq\mu,\quad i=1,\ldots,m,$
where $\alpha_{i}$ is the Lagrange multiplier. The decision function can be obtained by solving the above quadratic programming problem and is given by:
$f(x)=\operatorname{sgn}(w^{T}x+b)=\operatorname{sgn}\left(\sum_{i=1}^{m}\alpha_{i}y_{i}x_{i}^{T}x+b\right).$
A new sample is classified as $+1$ or $-1$ by utilizing the above decision function.
2.2 Twin Support Vector Machine
@@
where $w_{1},w_{2}\in\mathbb{R}^{n}$ and $b_{1},b_{2}\in\mathbb{R}$. Assume that the data points belonging to class $+1$ and class $-1$ are represented by the rows of matrices $A\in\mathbb{R}^{m_{1}\times n}$ and $B\in\mathbb{R}^{m_{2}\times n}$, respectively. Solving the following two quadratic programming problems yields the TSVM classifiers:
@@
where $C_{1}$, $C_{2}>0$ are penalty parameters, $\xi_{1}$, $\xi_{2}$ are slack vectors, and $e_{1}$, $e_{2}$ are vectors of ones of appropriate dimension. Using the KKT conditions, the Wolfe dual formulations of problems in (TSVM) are
$\mathop{\max}_{\alpha}$ $\quad-\frac{1}{2}\alpha^{T}G(H^{T}H)^{-1}G^{T}\alpha+e_{2}^{T}\alpha$
s.t. $\quad 0\leq\alpha\leq C_{1}e_{2},$
$\mathop{\max}_{\alpha^{*}}$ $\quad-\frac{1}{2}\alpha^{*T}H(G^{T}G)^{-1}H^{T}\alpha^{*}+e_{1}^{T}\alpha^{*}$
s.t. $\quad 0\leq\alpha^{*}\leq C_{2}e_{1},$
where $H=[A~{}~{}e_{1}]$, $G=[B~{}~{}e_{2}]$ and $\alpha$, $\alpha^{*}$ are the Lagrangian multipliers. After solving the above dual problems, the separating hyperplanes are found by
$\begin{bmatrix}w_{1}\\ b_{1}\end{bmatrix}=-(H^{T}H)^{-1}G^{T}\alpha,\qquad\mbox{and}\qquad\begin{bmatrix}w_{2}\\ b_{2}\end{bmatrix}=(G^{T}G)^{-1}H^{T}\alpha^{*}.$
In practice, a small perturbation of the identity matrix is often added to ensure the invertibility of the coefficient matrices above. The following decision rule is used to assign a new data point $x\in\mathbb{R}^{n}$ to class $k\in\{1,2\}$:
$\mbox{class}~{}k=\mbox{argmin}\dfrac{|w_{i}^{T}x+b_{i}|}{\|w_{i}\|^{2}},~{}~{}~{}i=1,2.$
2.3 Universum Twin Support Vector Machine
Universum data is defined as a set of unlabeled samples that do not belong to any specific class sinz2007analysis ; weston2006inference . This type of data can encode past knowledge by providing meaningful information within the same domain as the problem at hand. Incorporating Universum data has been shown to effectively improve learning performance in both classification and clustering tasks. By integrating Universum data into classical models, researchers have enhanced the computational efficiency and generalization ability of various existing methods moosaei2023universum ; moosaei2023sparse ; weston2006inference ; xiao2020new .
@@ 2.3 Universum Twin Support Vector Machine
where $C_{1},C_{2},$ and $C_{u}$ are positive penalty parameters, $\varepsilon\in(0,1)$ is the tolerance value for Universum class, $\xi_{1},\xi_{2},\psi_{1},$ and $\psi_{2}$ are measures of the violation of associated constraints. We may derive their dual problems by applying Lagrangian functions, as shown below.
@@
Once these vectors are obtained, the separating planes $w_{1}^{T}x+b_{1}=0$ and $w_{2}^{T}x+b_{2}=0,$ are known. A new data point $x\in\mathbb{R}^{n}$ is allocated to class $i\in\{+1,-1\}$ by a similar rule to the TSVM.
@@ 2.4 Least-Square Twin Support Vector Machine
Substituting the equality constraints of the above problems into associated objective functions, one obtains convex unconstrained optimization problems. Thus, to find their optimum solutions, we set their gradients with respect to $w_{1},b_{1},w_{2}$ and $b_{2}$ equal to 00. Consequently, we can find the two non-parallel hyperplanes by solving the following systems of linear equations
@@ 2.5 Least-Square Universum Twin Support Vector Machine
where $C_{1}$ and $C_{2}$ are positive penalty parameters, $\varepsilon\in(0,1)$ is the tolerance value for Universum class, $\xi_{1},\xi_{2},\psi_{1}$, and $\psi_{2}$ are measures of the violation of constraints associated. The unique solution to these problems is closed-form, which is advantageous for handling large-scale applications. By employing the same solution method as LS-TSVM, we can obtain the two non-parallel hyperplanes by solving the following system of linear equations.
@@ 2.6 Quadratic Twin Support Vector Machine
Linear models, while effective for linearly separable data, may face limitations when dealing with more complex datasets. To address this challenge, kernel methods are introduced. However, these methods have their drawbacks gao2019quadratic . In response, many quadratic kernel-free methods have been recently studied mousavi2019quadratic ; gao2021novel ; gao2019rescaled . The idea behind these models is analogous to their linear version except they capture quadratic surfaces. Below, we elaborate on the Quadratic Kernel-Free Twin Support Vector Machine (QTSVM) model proposed by Gao et al. gao2019quadratic .
Let $\mathcal{I}_{1}:=\{i\in\{1,\dots,n\}\ |\ y_{i}=1\}$ and $\mathcal{I}_{2}:=\{i\in\{1,\dots,n\}\ |\ y_{i}=-1\}$. Suppose we have two classes that include $|\mathcal{I}_{1}|$ and $|\mathcal{I}_{2}|$ points, respectively. Then, the formulation of the QTSVM model for binary classification becomes the following:
@@
It is important to note that the quadratic-based models are not in the standard form of quadratic programs. Therefore, we will introduce some notations and provide several definitions to facilitate their conversion into standard quadratic program forms. We shall be succinct here as these definitions are mainly borrowed from mousavi2019quadratic . For a square matrix $A=[a_{ij}]_{i=1,\dots,n;j=1,\dots,n}\in\mathbb{R}^{n\times n}$, its vectorization given by
@@
In case $A$ is symmetric, $\mbox{vec}(A)$ contains redundant information so that we often consider its half-vectorization given by:
@@
Given $n\in\mathbb{N}$, there exist a unique elimination matrix $L_{n}\in\mathbb{R}^{\frac{n(n+1)}{2}\times n^{2}}$ such that magnus1980elimination
@@
and further, this elimination matrix $L_{n}$ has full row rank magnus1980elimination . Conversely, for any $n\in\mathbb{N}$, there is a unique duplication matrix $D_{n}\in\mathbb{R}^{n^{2}\times\frac{n(n+1)}{2}}$ such that
@@
$s_{i}$ ::=:= $\frac{1}{2}\mbox{hvec}(x_{i}x_{i}^{T}),\qquad s_{j}:=\frac{1}{2}\mbox{hvec}(u_{j}u_{j}^{T}),\qquad r_{i}:=[s_{i};x_{i}],\quad r_{j}:=[s_{j};u_{j}],$
$w$ ::=:= $\mbox{hvec}(W),\quad z:=[w;b],\quad V:=\big{[}I_{\frac{n(n+1)}{2}}\,\,0_{\frac{n(n+1)}{2}\times n}\big{]},\qquad X_{i}:=I_{n}\otimes x_{i}^{T},$
$M_{i}$ ::=:= $X_{i}D_{n},\ \ H_{i}:=[M_{i}\,\,I_{n}],\,\ G:=2\sum_{i=1}^{m}H_{i}^{T}H_{i},\ \ X:=[x_{1}^{T};x_{2}^{T},\dots;x_{m}^{T};u_{1}^{T};u^{T}_{2};\dots;u_{r}^{T}].$
Consequently, for fixed $i=1,\dots,m$ and $j=1,\dots,r$, we get the following equations:
@@
Note that subindices $i$ and $j$ are consistently used to correspond with data points and Universum points throughout this paper, respectively. Using the aforementioned notations, let $z_{1}$ and $z_{2}$ be the decision variables constructed using $(W_{1},b_{1})$ and $(W_{2},b_{2})$, respectively. These variables are related to the quadratic surfaces for class 1 and class 2, respectively, which will be captured by twin models discussed later in the paper. Consequently, we can express the equivalent problem of QTSVM in the standard form of a quadratic minimization problem as follows:
@@
2.7 Least Square Quadratic Twin Support Vector Machine

### v2

@@ Abstract
Binary classification tasks with imbalanced classes pose significant challenges in machine learning. Traditional classifiers often struggle to accurately capture the characteristics of the minority class, resulting in biased models with subpar predictive performance. In this paper, we introduce a novel approach to tackle this issue by leveraging Universum points to support the minority class within quadratic twin support vector machine models. Unlike traditional classifiers, our models utilize quadratic surfaces instead of hyperplanes for binary classification, providing greater flexibility in modeling complex decision boundaries. By incorporating Universum points, our approach enhances classification accuracy and generalization performance on imbalanced datasets. We generated four artificial datasets to demonstrate the flexibility of the proposed methods. Additionally, we validated the effectiveness of our approach through empirical evaluations on benchmark datasets, demonstrating superior performance compared to conventional classifiers and existing methods for imbalanced classification.
@@ 1 Introduction
However, many real-world applications pose challenges, as they often involve data that is not linearly separable. Addressing this requires robust techniques capable of effectively handling nonlinear situations. One well-established approach is to leverage the concept of kernel methods, which relies on the existence of a nonlinear mapping. This mapping transforms the original data into a higher-dimensional, potentially infinite-dimensional feature space, where the data becomes linearly separable. Notably, the appeal of kernel methods lies in their ability to bypass the need for explicitly knowing this mapping. Instead, they focus on selecting appropriate kernel functions with advantageous properties. However, the challenge lies in determining the suitable kernel function, and fine-tuning the associated hyperparameters can be computationally demanding. As a result, there is a growing interest in practical methods that seek nonlinear classifiers directly in the original feature space.
Kernel-free models offer a direct approach for handling nonlinearly separable datasets without the need to map them to a higher-dimensional feature space. The quadratic surface support vector machine (QSSVM) dagher2008quadratic utilizes quadratic surfaces for class separation. Bai et al.bai2015quadratic introduced a kernel-free least squares QSSVM tailored for disease diagnosis. A kernel-free support vector machine, utilizing a double-well potential proposed in gao2021kernel , aims to achieve a specialized fourth-order polynomial separating surface. However, when pursuing nonlinear classifiers, the complexity of the optimization program significantly increases due to the involvement of numerous decision variables, leading to computational challenges and potential overfitting. To mitigate this, incorporating a surrogate that promotes sparsity among decision variables proves advantageous moosaei2023sparse ; mousavi2022quadratic ; mousavi2020survey . To achieve less expensive quadratic programs, gao2019quadratic discusses a least squares twin QSSVM, leveraging two quadratic surfaces for data representation.
@@
Gap in the literature.
Existing methods either (i) rely on kernel mappings with high computational cost, (ii) do not explicitly address class imbalance, or (iii) use quadratic surfaces without effectively incorporating additional prior knowledge (such as Universum points). As a result, there is a lack of models that simultaneously handle nonlinear decision boundaries, class imbalance, and incorporate prior knowledge in a computationally efficient framework.
Building on this gap, our methodology draws inspiration from the approach introduced in richhariya2020reduced , which incorporates Universum points to bolster the minority class in linearly separable datasets. Building upon this foundation, we extend its application to the framework of quadratic twin support vector machines (QTSVM). Leveraging the inherent flexibility of QTSVM, which employs quadratic surfaces instead of hyperplanes for binary classification, we aim to enhance the efficacy of Universum points in addressing class imbalance. This innovative integration enables us to capture intricate class relationships and adapt to complex data distributions more effectively. In addition to the hinge loss, we derive the least squares formulation, which adopts the $\ell_{2}$ norm to penalize slack variables and replaces inequality and equality constraints. To validate the effectiveness of our proposed methods, we conduct experiments on various datasets, including artificial and public benchmark datasets. These experiments demonstrate the efficacy and efficiency of our proposed models compared to other well-known SVM models.
In summary, the main contributions of this paper are as follows:
  • We introduce the Imbalanced Universum Quadratic Twin Support Vector Machine (Im-$\mathfrak{U}$-QTSVM), which leverages quadratic surfaces instead of hyperplanes to better capture complex decision boundaries in binary classification.
  • We incorporate Universum points to guide the minority class, improving generalization and reducing bias toward the majority class in imbalanced datasets.
  • We propose both hinge-loss and least squares formulations, enabling either standard QP solutions or closed-form solutions for computational efficiency.
  • We establish the theoretical properties and computational characteristics of the proposed models, including guaranteed optimality, uniqueness under mild conditions, sufficient conditions for nonzero solutions, and an analysis of computational complexity.
  • We demonstrate the effectiveness of the proposed methods through experiments on both artificial and benchmark datasets, showing superior performance compared to other existing approaches.
@@
$\mathbb{R}^{n}$ denotes the $n$-dimensional real vector space. For $x\in\mathbb{R}^{n}$, the notation $|x|$ signifies element-wise absolute value. $A^{T}$ represents the transpose of a matrix $A$ and $\|\cdot\|$ denotes the Euclidean norm. Consider a real-valued function $f$ defined on $\mathbb{R}^{n}$. The gradient of $f$ at a point $x$ is denoted by the $n$-dimensional column vector $\nabla f(x)$. Let $S_{n}$ denote the set of all $n\times n$ symmetric matrices. For a matrix $A\in S_{n}$, the notation $A\succ 0$ indicates that $A$ is positive definite.
@@ 2 Problem Statement and Related Works
2.1 Universum Twin Support Vector Machine
@@
where $w_{1},w_{2}\in\mathbb{R}^{n}$ and $b_{1},b_{2}\in\mathbb{R}$. Assume that the data points belonging to class +1+1 and class 1-1 are represented by the rows of matrices $A\in\mathbb{R}^{m_{1}\times n}$ and $B\in\mathbb{R}^{m_{2}\times n}$, respectively. Solving the following two quadratic programming problems yields the TSVM classifiers:
@@
where $C_{1}$, $C_{2}>0$ are penalty parameters, $\xi_{1}$, $\xi_{2}$ are slack vectors, and $e_{1}$, $e_{2}$ are vectors of ones of appropriate dimension.
Universum data is defined as a set of unlabeled samples that do not belong to any specific class sinz2007analysis ; weston2006inference . This type of data can encode past knowledge by providing meaningful information within the same domain as the problem at hand. Incorporating Universum data has been shown to effectively improve learning performance in both classification and clustering tasks. By integrating Universum data into classical models, researchers have enhanced the computational efficiency and generalization ability of various existing methods moosaei2023universum ; moosaei2023sparse ; weston2006inference ; xiao2021new .
@@
where $C_{1},C_{2},$ and $C_{u}$ are positive penalty parameters, $\varepsilon\in(0,1)$ is the tolerance value for the Universum class, and $\xi_{1},\xi_{2},\psi_{1},$ and $\psi_{2}$ are measures of the violation of the associated constraints. We may derive their dual problems by applying Lagrangian functions, as shown below.
@@
Once these vectors are obtained, the separating planes $w_{1}^{T}x+b_{1}=0$ and $w_{2}^{T}x+b_{2}=0,$ are known. A new data point $x\in\mathbb{R}^{n}$ is allocated to class $i\in\{+1,-1\}$ by a rule similar to the TSVM.
@@ 2.2 Least-Square Twin Support Vector Machine
Substituting the equality constraints of the above problems into the associated objective functions, one obtains convex unconstrained optimization problems. Thus, to find their optimum solutions, we set their gradients with respect to $w_{1},b_{1},w_{2}$ and $b_{2}$ equal to 0. Consequently, we can find the two non-parallel hyperplanes by solving the following systems of linear equations
@@ 2.3 Least-Square Universum Twin Support Vector Machine
where $C_{1}$ and $C_{2}$ are positive penalty parameters, $\varepsilon\in(0,1)$ is the tolerance value for the Universum class, and $\xi_{1},\xi_{2},\psi_{1}$ and $\psi_{2}$ are measures of the violation of associated constraints. The unique solution to these problems is closed-form, which is advantageous for handling large-scale applications. By employing the same solution method as LS-TSVM, we can obtain the two non-parallel hyperplanes by solving the following system of linear equations.
@@ 2.4 Quadratic Twin Support Vector Machine
Linear models, while effective for linearly separable data, may face limitations when dealing with more complex datasets. To address this challenge, kernel methods have been introduced. However, these methods have their drawbacks gao2019quadratic . In response, many quadratic kernel-free methods have been studied recently mousavi2022quadratic ; gao2021novel ; gao2019rescaled . The idea behind these models is analogous to that of their linear version, except they capture quadratic surfaces. Below, we elaborate on the Quadratic Kernel-Free Twin Support Vector Machine (QTSVM) model proposed by Gao et al. gao2019quadratic .
Suppose we are given a dataset of $n$ points $\{(x_{i},y_{i})\}_{i=1}^{n}$, where $x_{i}$ are the feature vectors and $y_{i}\in\{-1,+1\}$ are the binary class labels. We can partition the indices of the data points based on their labels as follows: $\mathcal{I}_{1}:=\{i\in\{1,\dots,n\}\ |\ y_{i}=1\}$ and $\mathcal{I}_{2}:=\{i\in\{1,\dots,n\}\ |\ y_{i}=-1\}$. Suppose we have two classes that include $|\mathcal{I}_{1}|$ and $|\mathcal{I}_{2}|$ points, respectively. Then, the formulation of the QTSVM model for binary classification becomes the following:
@@
It is important to note that the quadratic-based models are not in the standard form of quadratic programs. Therefore, we will introduce some notations and provide several definitions to facilitate their conversion into standard quadratic program forms. We shall be succinct here, as these definitions are mainly borrowed from mousavi2019quadratic. For a square matrix $A=[a_{ij}]_{i=1,\dots,n;j=1,\dots,n}\in\mathbb{R}^{n\times n}$, its vectorization is given by
@@
In case $A$ is symmetric, $\mbox{vec}(A)$ contains redundant information, so we often consider its half-vectorization given by:
@@
Given $n\in\mathbb{N}$, there exists a unique elimination matrix $L_{n}\in\mathbb{R}^{\frac{n(n+1)}{2}\times n^{2}}$ such that magnus1980elimination
@@
and furthermore, this elimination matrix $L_{n}$ has full row rank magnus1980elimination . Conversely, for any $n\in\mathbb{N}$, there is a unique duplication matrix $D_{n}\in\mathbb{R}^{n^{2}\times\frac{n(n+1)}{2}}$ such that
@@
$s_{i}$ $:=$ $\frac{1}{2}\mbox{hvec}(x_{i}x_{i}^{T}),\qquad s_{j}:=\frac{1}{2}\mbox{hvec}(u_{j}u_{j}^{T}),\qquad r_{i}:=[s_{i};x_{i}],\quad r_{j}:=[s_{j};u_{j}],$
$w$ $:=$ $\mbox{hvec}(W),\quad z:=[w;b],\quad V:=\big[I_{\frac{n(n+1)}{2}}\,\,0_{\frac{n(n+1)}{2}\times n}\big],\qquad X_{i}:=I_{n}\otimes x_{i}^{T},$
$M_{i}$ $:=$ $X_{i}D_{n},\ \ H_{i}:=[M_{i}\,\,I_{n}],\,\ G:=2\sum_{i=1}^{m}H_{i}^{T}H_{i},\ \ X:=[x_{1}^{T};x_{2}^{T},\dots;x_{m}^{T};u_{1}^{T};u^{T}_{2};\dots;u_{r}^{T}].$
Consequently, for fixed $i=1,\dots,m$ and $j=1,\dots,r$, we obtain the following equations:
@@
Note that subindices $i$ and $j$ are consistently used to correspond with data points and Universum points throughout this paper, respectively. Using the aforementioned notations, let $z_{1}$ and $z_{2}$ be the decision variables constructed using $(W_{1},b_{1})$ and $(W_{2},b_{2})$, respectively. These variables are related to the quadratic surfaces for class 1 and class 2, respectively, which will be captured by the twin models discussed later in the paper. Consequently, we can express the equivalent problem of QTSVM in the standard form of a quadratic minimization problem as follows:
@@
2.5 Least Squares Quadratic Twin Support Vector Machine